OpenStudy (anonymous):

How much work can be extracted from 5.00 × 10^4 J of heat when 2.00 × 10^4 J of heat is exhausted? A. - 7.00 × 10^4 J B. - 3.00 × 10^4 J C. 3.00 × 10^4 J D. 7.00 × 10^4 J ***thank you:)

OpenStudy (michele_laino):

I think that we have to apply the first principle of thermodynamics

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so, we can write this: \[Q = L + \Delta U\]

OpenStudy (michele_laino):

where Q= 2*10^4 Joules, and \Delat U = 5*10^4 Joules

OpenStudy (michele_laino):

oops..\Delta*

OpenStudy (anonymous):

okay!! so we get 70000? so our solution is choice D?

OpenStudy (michele_laino):

no,we have: \[2 \times {10^4} = L + 5 \times {10^4}\]

OpenStudy (anonymous):

ohh oops!! so we get 20000 = L + 50000 L = - 30000 ? so we solution is choice B?

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

yay:) thank you!! :)

OpenStudy (michele_laino):

thank you!! :D